# A Practitioner’s Guide to Kolmogorov–Arnold Networks

### v1

@@ 2 MLP Basics
3 KAT and KAN
Kolmogorov–Arnold Theorem (KAT). A landmark result by Kolmogorov [122, 123] established that every continuous multivariate function can be expressed as a superposition of continuous univariate functions. Specifically, for any $f:[0,1]^{n}\to\mathbb{R}$,
$f(x_{1},\ldots,x_{n})=\sum_{q=0}^{2n}\underbrace{\Phi_{q}\!\Bigg(\,\sum_{p=1}^{n}\phi_{q,p}(x_{p})\Bigg)}_{\text{outer nonlinear map}},$ (4)
where each inner map $\phi_{q,p}:[0,1]\to\mathbb{R}$ and each outer map $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ is continuous. A key feature is that the inner functions $\{\phi_{q,p}\}$ can be chosen universal, i.e., independent of the specific target $f$, while the outer functions $\{\Phi_{q}\}$ encode dependence on $f$ itself [124, 125].
In following, we use $C(\cdot)$ to denote the space of continuous functions. The class ${\rm Lip}^{\alpha}$ refers to Hölder/Lipschitz functions of order $\alpha\in(0,1]$: there exists $L>0$ such that
$|u(x)-u(y)|\leq L\,|x-y|^{\alpha}.$
The special case $\alpha=1$ corresponds to the standard Lipschitz condition. Fixed positive weights are written $\lambda_{p}$, shifts as $\eta_{q}$, offsets as $c_{q}$, and the notation $\nearrow$ indicates strictly increasing functions.
Simplified constructive variants. Since Kolmogorov’s original proof used highly irregular functions, later refinements proposed more structured forms:
  • Lorentz [126]. The representation can be simplified by using a single outer function $\Phi$ applied to weighted sums of smoother inner functions:
    $f(x_{1},\ldots,x_{n})=\sum_{q=0}^{2n}\Phi\!\Bigg(\sum_{p=1}^{n}\lambda_{p}\,\psi_{q}(x_{p})\Bigg),\quad\psi_{q}\in{\rm Lip}^{\alpha},\;\;\psi_{q}\nearrow.$
  • Sprecher [129]. The inner layer can be compressed further by reusing a single inner function $\phi$, shifted and offset differently across terms:
    $f(x_{1},\ldots,x_{n})=\sum_{q=0}^{2n}\Phi\!\Bigg(\sum_{p=1}^{n}\lambda_{p}\,\phi(x_{p}+\eta_{q})+q\Bigg).$ (5)
  • Ostrand [130]. Extended the two-layer superposition representation beyond cubes $0,1]^{n}$ to general compact metric spaces, broadening its applicability.
Note that the exact inner functions constructed in KAT are often pathological [131]. Modern refinements instead approximate both $\phi$ and $\Phi$ using smooth, learnable parameterizations—an idea that directly motivates KANs.
From KAT to KAN. KANs [27] translate the KAT blueprint into a trainable neural layer by learning the univariate maps and summing them:
$x^{(\ell+1)}_{q}=\sum_{p=1}^{P}\underbrace{\varphi^{(\ell)}_{q,p}\!\big(x^{(\ell)}_{p}\big)}_{\begin{subarray}{c}\text{edge-wise}\\ \text{1D transform}\end{subarray}},\quad q=1,\dots,Q,$ (6)
where $x^{(\ell)}_{p}$ is the $p$-th coordinate of the input at layer $\ell$, $P$ is the input dimension, and $Q$ is the output dimension. Compared with MLPs (which first form a linear combination and then apply a pointwise activation), KANs apply per-coordinate univariate transforms first and then aggregate additively.
Learning the 1D maps. A common parameterization is a basis expansion, e.g., with B-spline functions:
$\varphi^{(\ell)}_{q,p}(x)=\sum_{k=1}^{K}\underbrace{c^{(\ell)}_{q,p,k}}_{\begin{subarray}{c}\text{learnable}\\ \text{coefficient}\end{subarray}}\underbrace{B_{k}(x)}_{\begin{subarray}{c}\text{basis}\\ \text{function}\end{subarray}},$ (7)
where $\{B_{k}\}_{k=1}^{K}$ are spline basis functions, and their locality/smoothness yields compact and interpretable representations of $\varphi$. Other bases (Chebyshev, Jacobi, Fourier, radial basis functions, or ReLU powers) can be substituted directly; the structure in (6) remains unchanged, only the parameterization of the univariate maps $\phi$ or $\varphi$ differs.
Remark. Earlier depth-2, width-(2n+1)(2n{+}1) realizations of KAT were difficult to optimize in practice [132, 133, 134, 135, 136]; modern KANs address this via differentiable basis functions and end-to-end training [27].
@@ 6 Basis Functions
This section reviews commonly used bases in KANs: B-splines, Chebyshev and Jacobi polynomials, ReLU compositions, Fourier series, Gaussian kernels, wavelets, finite-basis partitions, and Sinc functions.
We begin with an overview (Table 5) to orient the reader. Subsequent subsections expand on each family, beginning with B-splines.
@@ 6.1 B-spline
A global tanh is applied after each hidden layer (but not the output) to keep activations within the spline domain. KAN 2 further supports post–training grid refinement, which increases knot resolution without reinitializing parameters, thereby improving fidelity with minimal overhead. Further discussion on efficiency and accuracy improvements of B-spline KANs can be found in Sec. 7 and Sec. 8, respectively.
@@ 6.2 Chebyshev Polynomials
consistent with the generic KAN layer formulation (6).
@@
Normalization and Stability. Figure 2 illustrates how per-layer $\tanh$ normalization stabilizes Chebyshev activations. Panel (a) shows the standard basis $T_{k}(x)$ on [1,1][-1,1] for degree K=8K{=}8, where extrema occur at $T_{k}(\pm 1)=\pm 1$, forcing steep slopes near the boundaries. Panel (b) instead evaluates $T_{k}(\tanh x)$, compressing the effective range to $\tanh(\pm 1)\approx\pm 0.762$, which moderates endpoint slopes while preserving interior structure. This compression prevents saturation at $\pm 1$ and keeps features in a well-conditioned range—critical for deep stacks of Chebyshev layers.
@@
Grid-Averaged Variant. Following Toscano et al. [44], a practical enhancement introduces a small learnable grid before evaluating the Chebyshev expansion. For a univariate input $x$,
@@
where $m$ denotes the number of centers, $b_{i}$ are uniformly spaced biases within $\beta_{\min},\beta_{\max}]$ (e.g., 0.1-0.1 to $0.1$), and $w_{i}$ control local slope. The coefficients $c^{(\ell)}_{q,p,k}$ can follow the stable initialization proposed in [169]. This averaged-$\tanh$ grid constrains $g(x)\in(-1,1)$, introduces mild spatial warping, improves numerical conditioning, and—as reported in [44]—supports stable training even with higher polynomial degrees.

### v2

@@ 2 MLP Basics
3 KST Toward Neural Networks
Superpositions and Hilbert’s Question. A superposition is any function obtained by repeatedly composing simpler functions. Lorentz [82] illustrates this idea with the example
$F(x,y,z)=f\!\bigl(g(x,y),\,h(x,\,k(y,z))\bigr),$
for scalar inputs $x,y,z\in\mathbb{R}$, showing how a multivariate mapping can be built from bivariate ones. Even the elementary identity
$x_{1}+\cdots+x_{n}=x_{1}+\bigl(x_{2}+\cdots+(x_{n-1}+x_{n})\cdots\bigr)$
is a superposition of binary additions.
Hilbert’s 13th Problem [76], posed at the 1900 International Congress of Mathematicians, asked whether every continuous function of three variables
$f:[0,1]^{3}\to\mathbb{R}$
can be represented as a finite superposition of continuous bivariate functions. Hilbert conjectured that such a representation was impossible. However, the later work of Kolmogorov and Arnol’d overturned this view, showing that every continuous function of $s$ variables can be expressed using only sums and superpositions of univariate functions—a result that fundamentally reshaped the theory of functional representation.
Kolmogorov and Arnol’d: resolving the 13th Problem. Kolmogorov’s 1956 note [77] introduced the idea that multivariate functions might be representable via functions of fewer variables. Arnol’d, his student, confirmed this for the three–variable case in 1957 [78], giving an explicit decomposition of a general $f(x,y,z)$ into superpositions of continuous bivariate functions. Later that year, Kolmogorov published his full and elementary solution [79]: any continuous function of $n>2$ variables can be written using only univariate functions and addition. Thus Hilbert’s conjecture was disproved.
The Kolmogorov Superposition Theorem (KST). Let $E_{1}=[0,1]$ and $E_{n}=[0,1]^{n}$. Kolmogorov proved that for every $n>2$ there exist continuous universal inner functions
$\phi_{q,p}:E_{1}\to\mathbb{R},\qquad p=1,\dots,n,\;\;q=0,\dots,2n,$
independent of the target function $f$, such that every $f\in C(E_{n})$ admits the representation
$f(x_{1},\ldots,x_{n})=\sum_{q=0}^{2n}\Phi_{q}\!\Bigg(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\Bigg),\qquad(x_{1},\ldots,x_{n})\in E_{n},$ (4)
where each $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ is continuous and depends on $f$, while the inner functions $\phi_{q,p}$ are universal. KST shows that multivariate continuity can be synthesized entirely from univariate nonlinearities, a structural insight that later motivated two–layer interpretations and neural–network analogies.
Proof strategies and later refinements. Two major approaches emerged in the decades following Kolmogorov’s work:
(i) Baire–category proofs: Kahane [80] and Hedberg [81] showed that for generic tuples of continuous increasing functions, the superposition operator is surjective. Their arguments reveal that KST is not only true but abundant: for almost all choices of inner functions, an appropriate outer function exists.
(ii) Constructive proofs: Lorentz [82], Sprecher [83], Arnol’d [84], and others developed explicit (though highly involved) constructions. These proofs follow a common pattern: partition $E_{n}$ into subcubes, map each subcube into a disjoint real interval via the inner functions, and build the outer functions iteratively to approximate the target $f$. Kolmogorov’s original paper left several analytic details unstated; these were filled in by Arnol’d [84] and later authors, making the constructive picture complete.
This classical theory forms the starting point for the refinements described next: reductions in the number of unique functions, improvements in regularity, and the development of computationally usable versions that ultimately connect the superposition viewpoint to modern KAN architectures.
Refinements to KST. Kolmogorov’s constructive formulation (4) established the universality of the inner functions $\phi_{q,p}$, but it did so at considerable structural cost: the representation requires $2n^{2}{+}n$ distinct inner functions and 2n+12n{+}1 outer functions. While theoretically elegant, this level of redundancy makes the original form far too complex for computation or numerical approximation. The decades following Kolmogorov’s work therefore focused on two goals: (i) reducing the number of unique functions needed, and (ii) improving their regularity. These efforts produced the modern, compressed form of KST that underlies computational constructions and later neural interpretations.
(i) Reducing the number of functions. Sternfeld [85] proved that the number of summands $2n+1$ in the Kolmogorov superposition formula is optimal and cannot be reduced. However, his functional–analytic framework shows that these 2n+12n{+}1 summands do not require 2n+12n{+}1 different inner or outer functions: the same small collection of functions can be reused repeatedly. Thus the number of distinct functions can be much smaller, even though the number of summands is fixed.
Lorentz’ reduction of outer functions. Lorentz’s 1962 exposition [82] clarified Kolmogorov’s argument and showed that the 2n+12n{+}1 outer functions $\Phi_{q}$ can be replaced by a single continuous function $\Phi$. In the bivariate case ($s=2$), with inputs $(x_{1},x_{2})\in[0,1]^{2}$, he derived the canonical representation
$f(x_{1},x_{2})=\sum_{q=1}^{5}\Phi\!\bigl(\phi_{1,q}(x_{1})+\phi_{2,q}(x_{2})\bigr),$
where the ten inner functions $\phi_{1,q}$ and $\phi_{2,q}$ are universal, but the outer function $\Phi$ depends on $f$. More generally, for $s$ variables he obtained
$f(x_{1},\ldots,x_{s})=\sum_{q=1}^{2s+1}\Phi\!\left(\sum_{p=1}^{s}\phi_{p,q}(x_{p})\right),$ (5)
demonstrating that the dependence on $f$ enters only through a single continuous outer function $\Phi$.
Sprecher’s reduction of inner functions. Sprecher [83] achieved the corresponding simplification for the inner layer. He showed that the entire family $\{\phi_{q,p}\}$ can be generated from a single monotone universal function $\psi:E_{1}\to\mathbb{R}$, together with fixed positive weights $\{\lambda_{p}\}_{p=1}^{n}$ and uniform shifts $x\mapsto x+\varepsilon q$, where $\varepsilon>0$ is a small constant. In modern notation,
$\phi_{q,p}(x)=\lambda_{p}\,\psi(x+\varepsilon q),$ (6)
where the weights $\lambda_{p}>0$ depend only on the dimension and satisfy a rational independence condition. Sprecher also proved that $\psi$ may be chosen with an explicit Hölder–Lipschitz exponent.
By combining Lorentz’ outer-function reduction and Sprecher’s inner-function compression, one obtains the modern compressed version of the superposition theorem:
$f(x_{1},\ldots,x_{n})=\sum_{q=0}^{2n}\Phi\!\Bigg(\sum_{p=1}^{n}\lambda_{p}\,\psi(x_{p}+\varepsilon q)+\delta q\Bigg),$ (7)
where $\psi:E_{1}\to\mathbb{R}$ is a single universal inner function, $\Phi:\mathbb{R}\to\mathbb{R}$ is a continuous outer function depending on $f$, the constants $\{\lambda_{p}\}$ are positive and rationally independent, $\varepsilon>0$ is a uniform shift, and $\delta\in\mathbb{R}$ is an offset introduced for separation of summands.
(ii) Improving regularity. The inner functions in Kolmogorov’s original proof are highly irregular—Hölder but non-differentiable. A function $u:E_{1}\to\mathbb{R}$ is Hölder continuous with exponent $\alpha\in(0,1]$ if
$|u(x)-u(y)|\leq L|x-y|^{\alpha},\qquad x,y\in E_{1},$
for some $L>0$. The case $\alpha=1$ corresponds to Lipschitz continuity.
Vitushkin and Henkin [86, 87, 88] proved that no universal inner family can possess $C^{1}$ (or $C^{s}$ with $s>1$) smoothness. Within the theoretical development of KST itself, Fridman [92] constructed universal inner maps in $\mathrm{Lip}^{1}$, the maximal regularity consistent with these impossibility results. Sprecher [93] further demonstrated that his compressed formulation (6) can also be realized with Lipschitz inner functions. More recently, Actor and Knepley introduced the first explicit algorithm aimed at computing Lipschitz universal inner functions, refining the earlier Hölder-type constructions of Köppen and Braun–Griebel [94, 95].
Computational KST. Algorithmic approaches fall into three categories:
  • Exact constructive KST: compute the universal inner map and the outer functions exactly.
  • Approximate KST (outer-only): fix the universal inner map and approximate or learn only the outer function.
  • Approximate KST (inner–and–outer): approximate both inner and outer maps within a superposition-type architecture.
Exact constructions. Sprecher’s digit-based inner function and outer correction schemes were made fully computable and later corrected by Köppen and rigorously validated by Braun–Griebel, who established continuity and monotonicity of the corrected recursion [96, 97, 98, 99]. Constructive and guaranteed versions include Nakamura et al., who provide algorithmic constructions of the universal inner map $\psi$ and outer map $\chi$ with certified error bounds [101].
Approximate KST (outer-only; inner fixed). This line of work preserves the universal inner map and reduces learning to a one-dimensional outer problem.
  • De Figueiredo (1980): fixes Sprecher’s universal inner function $\psi$ and learns the outer function $g$ using a Chebyshev basis—the first data-driven outer-fitting scheme in a KST layout [104].
  • Frisch et al. (1989): retain Lorentz’s universal inner maps and iteratively refine a single outer function $g$ [111].
  • Braun–Griebel inner + RKHS outer (2009): uses the corrected universal inner function $\psi$ and replaces constructive outer windows by spline/Fourier RKHS approximants [100].
  • Lai–Shen (2025): again fix the Lorentz–Sprecher inner maps and reconstruct the outer function using linear B–splines with $O(n^{-2})$ accuracy [265, 266].
Outer-only methods preserve inner universality and thus avoid the curse of dimensionality: all problem dependence flows through a univariate outer approximation.
Approximate KST (inner–and–outer; controversial). These approaches approximate or entirely replace the universal inner functions, thereby departing from the exact Kolmogorov formulation.
  • Kurková (1991–1992): replaces the universal inner staircase by sigmoidal staircase approximants and learns the outer coefficients within a two-hidden-layer sigmoidal architecture [90, 91].
  • Nees (1994, 1996): constructs explicit recursive piecewise-linear approximations of the inner maps $\phi_{q}^{k}$ together with data-driven outer corrections $g_{r}$ that achieve geometric error decay [102, 103].
  • Spline-based architectures: The KSN model of Igelnik and Parikh (2003) [106] uses cubic splines for both inner and outer functions, while Coppejans (2005) [107] introduces monotone inner splines for statistical identifiability. Related spline-based superpositional models, including ExSpliNet [108], KASAM [109], and the Urysohn-operator framework of Polar and Poluektov [110], adopt additive or hierarchical structures reminiscent of KST but do not enforce universality of the inner layer. These methods yield practical spline architectures but are not constructive realizations of the Kolmogorov superposition theorem.
Caveat. Exact KST requires a universal (target-independent) inner function; approximating or smoothing this function breaks universality and may violate separation. Classical results of Vitushkin and Henkin, and the modern analysis of Actor–Knepley, show that no $C^{1}$ or overly smooth inner functions can satisfy the Kolmogorov separation property [86, 88, 87, 95]. Thus inner approximations are meaningful only in the approximation-theoretic sense, not as exact KST.
Neural interpretations. De Figueiredo (1980) [104] was the first to view the Kolmogorov–Sprecher representation as a computational architecture, expressing it in block-diagram form and proposing a data-driven procedure for learning the outer function $g$ using Chebyshev bases. Hecht–Nielsen (1987) [105] interprets Sprecher’s representation as a three-stage mapping network with a fixed universal hidden layer and learned outer functions. Kurková’s approximation framework (1991–1992) [90, 91] provides compatibility with standard trainable neural networks by replacing the inner staircase with smooth sigmoidal combinations while learning the outer layer. Spline-based models such as KSN (2003) [106] and the shape-constrained variant of Coppejans (2005) [107] train both inner and outer functions within spline families, yielding practical but non-KST architectures. See Table 4 for a chronological overview of these developments, tracing the progression from the original KST theory to its computational and neural interpretations.
Table 4: Development of Kolmogorov’s Superposition Theorem towards Kolmogorov-inspired networks (1900–2024).
Year Author(s) Key Contribution Focus on
1900 Hilbert [76] Poses Hilbert’s 13th problem Theory
1956 Kolmogorov [77] Preliminary idea of superpositions; first hint toward the theorem. Theory
1957 Arnol’d [78] First explicit 3-variable construction with 9 terms; provides concrete counterexample to Hilbert 13. Theory
1957 Kolmogorov [79] Full Kolmogorov Superposition Theorem; first general $n$-D proof, Theory
1958 Arnol’d [84] Supplies missing lemmas; first to complete Kolmogorov’s proof. Theory
1962 Lorentz [82] Gives simplified canonical form with one outer function. Theory
1965 Sprecher [83] First single universal inner function; reduces many $\phi_{p,q}$ to one $\psi$ with shifts; major structural simplification over Kolmogorov. Theory
1967 Fridman [92] Shows universal inner functions may be taken $\mathrm{Lip}(1)$; first optimal smoothness result; strengthens Sprecher–Kolmogorov scheme. Theory
1980 de Figueiredo [104] First network-like interpretation; introduces block diagram + data-driven outer fitting; operationalizes Sprecher’s form. Theory
1987 Hecht–Nielsen [105] First explicit neural mapping theorem; views KST as two-hidden-layer network. Network
1989 Girosi–Poggio [89] First critical analysis of learnability; shows inner function must be non-smooth + outer non-parametric. Theory
1989 Frisch et al. [111] First computational implementation of Lorentz form; iterative learning of single outer function. Network
1991 Kurková [90] First approximation-theoretic reinterpretation; ties network size to modulus of continuity. Network
1992 Kurková [91] Two-hidden-layer sigmoidal approximants; universal inner parameters. Both
1993 Sprecher [112] Single universal $\psi$ for all $n$. Both
1993 Nakamura et al. [101] First fully constructive + guaranteed-accuracy version. Theory
1994 Nees [102] First use of spline (piecewise-linear) inner maps, introducing geometric error decay and a fully constructive, algorithmic approximation of KST with explicit error bounds; Both
1996 Sprecher [96] First executable version of $\psi$ with separation guarantees. Both
1997 Sprecher [97] Explicit algorithm for outer functions. Both
2002 Köppen [98] Corrected continuous monotone inner function of Sprecher [96]; first training-ready inner map. Network
2003 Igelnik–Parikh [106] First fully trainable Kolmogorov Spline Network; replaces fractal inner functions with smooth cubic splines and provides learnable inner and outer maps. Network
2009 Braun–Griebel [99] First correct constructive KST; fixes Sprecher’s errors [96]; modern constructive foundation. Theory
2019 Actor–Knepley [95] Proves $C^{1}$ inner functions impossible; first rigorous smoothness obstruction. Theory
2024 Liu et al. [26] First deep network architecture inspired by the Kolmogorov–Arnold representation (KAN). Network
2024: Liu et al. — Kolmogorov–Arnold Network
Liu et al. [26] introduce the Kolmogorov–Arnold Network, a neural architecture explicitly motivated by the structure of the KST (4), though not intended as a mathematical realization of it. The authors emphasize that KAN is not an exact implementation of the Kolmogorov Superposition Theorem, but rather a computational architecture “inspired by” the Kolmogorov–Arnold representation.
Because the Kolmogorov Superposition Theorem prescribes only a two-layer structure (universal inner functions $\phi_{q,p}$ followed by outer maps $\Phi_{q}$), Liu et al. generalize the idea by redefining the basic computational block. A KAN layer is introduced as a matrix of univariate functions,
$\bm{\Phi}_{\ell}=\{\phi_{\ell,j,i}\},\qquad i=1,\ldots,n_{\ell},\;\;j=1,\ldots,n_{\ell+1},$
where $n_{\ell}$ and $n_{\ell+1}$ denote the widths of layers $\ell$ and $\ell+1$, respectively, and where each edge $(i\!\to\!j)$ carries a trainable B-spline activation $\phi_{\ell,j,i}:\mathbb{R}\to\mathbb{R}$. Within a layer, the forward computation is purely Kolmogorov-like aggregation:
$x_{\ell+1,j}=\sum_{i=1}^{n_{\ell}}\phi_{\ell,j,i}\!\bigl(x_{\ell,i}\bigr),$ (8)
so each layer performs only “inner-type” operations: univariate nonlinearities followed by summation.
In the deep setting, the outputs of one collection of inner-type functions become the inputs to the next. No separate outer function $\Phi_{q}$ is introduced: the final layer plays the role of an outer map only in the trivial sense that it produces the scalar output. Thus a deep KAN is simply the composition of its inner-type layers,
$\mathrm{KAN}(\mathbf{x})=\bigl(\bm{\Phi}_{L-1}\circ\cdots\circ\bm{\Phi}_{0}\bigr)(\mathbf{x}),$ (9)
with “outer-like” effects distributed across all layers rather than concentrated in a single KST-style stage.
Why KANs should not be viewed as exact realizations of the Kolmogorov Superposition Theorem. Although Liu et al. describe KANs as inspired by the Kolmogorov superposition theorem, it is helpful to distinguish KST and KAN as follows:
  • Inner functions are learned, not universal. In KST the inner maps are universal and fixed for all target functions [79, 83], whereas in KANs every edge function is trained from data. This points to a structural analogy rather than an exact correspondence.
  • Regularity properties differ. Classical KST inner functions are only Hölder or Lipschitz continuous [86, 88, 92], whereas KANs use smooth B–splines. This indicates that KANs operate within an approximation framework rather than reproducing the exact KST construction.
  • Separation constraints are not enforced. Constructive proofs of KST rely on fixed shifts and integrally independent weights to separate subcube images [83, 82]. KANs do not impose these geometric constraints, so they do not replicate the mechanism underlying the theorem.
  • Width and depth do not follow KST. Sternfeld’s optimal width 2n+12n{+}1 applies to the shallow superposition setting [85]. KANs allow arbitrary widths and depths, reflecting modern deep-learning practice rather than KST optimality.
  • No final one-dimensional outer function. KST reduces approximation to a single univariate outer map [82, 83]. Deep KANs distribute nonlinearity across multiple layers and do not preserve this one-dimensional reduction.
@@ 6 Basis Functions
This section reviews commonly used bases in KANs: B-splines, Chebyshev and Jacobi polynomials, ReLU compositions, Fourier series, Gaussian kernels, wavelets, finite-basis partitions, and Sinc functions. We begin with an overview (Table 6) to orient the reader. Subsequent subsections expand on each family, beginning with B-splines.
@@ 6.1 B-spline
A global tanh is applied after each hidden layer (but not the output) to keep activations within the spline domain. KAN 2 further supports post–training grid refinement, which increases knot resolution without reinitializing parameters, thereby improving fidelity with minimal overhead.
Convergence implications. Recent theory gives spline-KANs some of the strongest guarantees in the KAN family. Wang [53] proves accelerated Sobolev approximation rates; Kratsios et al. [271] show optimal Besov performance on Lipschitz and even fractal domains; and Liu et al. [140] establish minimax-optimal regression rates together with optimal spline–knot scaling. These results place B-splines on the firmest theoretical footing among KAN bases. (See Sec. 10.)
Spectral bias and NTK conditioning. Spline grids control frequency reach: finer grids reduce spectral bias but may worsen conditioning [53, 39]. Initialization also matters—Rigas et al. [276] [^14] https://github.com/srigas/KAN_Initialization_Schemes show that Glorot-style and power-law initializations yield more stable NTKs and faster early convergence. (See Secs. 10.2 and 10.3.)
@@ 6.2 Chebyshev Polynomials
consistent with the generic KAN layer formulation.
@@
Normalization and Stability. Figure 2 illustrates how per-layer $\tanh$ normalization stabilizes Chebyshev activations. Panel (a) shows the standard basis $T_{k}(x)$ on [1,1][-1,1] for degree K=8K{=}8, where extrema occur at $T_{k}(\pm 1)=\pm 1$, forcing steep slopes near the boundaries. Panel (b) instead evaluates $T_{k}(\tanh x)$, compressing the effective range to $\tanh(\pm 1)\approx\pm 0.762$, which moderates endpoint slopes while preserving interior structure. This compression prevents saturation at $\pm 1$ and keeps features in a well-conditioned range, critical for deep stacks of Chebyshev layers.
@@
Grid-Averaged Variant. Following Toscano et al. [52], a practical enhancement introduces grid before evaluating the Chebyshev expansion. For a univariate input $x$,
@@
where $m$ denotes the number of centers, $b_{i}$ are uniformly spaced biases within $\beta_{\min},\beta_{\max}]$ (e.g., 0.1-0.1 to $0.1$), and $w_{i}$ control local slope. The coefficients $c^{(\ell)}_{q,p,k}$ can follow the stable initialization proposed in [196]. This averaged-$\tanh$ grid constrains $g(x)\in(-1,1)$, introduces mild spatial warping, improves numerical conditioning, and as reported in [52], supports stable training even with higher polynomial degrees.
@@
Depth Stabilization. Deep Chebyshev KANs often become unstable. Rigas et al. [262] [^16] https://github.com/srigas/RGA-KANs mitigate this with a basis-agnostic Glorot-style initialization that preserves activation variance, and a Residual–Gated Adaptive (RGA) KAN block that stabilizes deep cPIKANs on PDEs. One-liner: variance-preserving init + residual gating enable deep Chebyshev KANs to train without divergence.